# Understanding and Leveraging the Expert Specialization of Context Faithfulness in Mixture-of-Experts LLMs

### v1

@@ 6 Conclusion
Looking forward, we see several exciting directions: combining mechanistic interpretability techniques such as SAE Kang et al. (2025) to further unravel MoE experts; extending our methods to other forms of specialization such as reflection Li et al. (2025) and reasoning Liu et al. (2025). We hope this work inspires further research into aspect-specific expert discovery and optimization.
@@ Appendix A Implementation Details
We fine-tune the model for 1 epoch on SQuAD, NQ, and HotpotQA datasets, and 3 epochs on NQ-Swap and ConfiQA datasets. All experiments use the AdamW Loshchilov and Hutter (2019) optimizer with a learning rate of 5e-4, a warmup ratio of 0.1, and cosine learning rate decay. The batch size is set to 8, and the maximum sequence length is 300. All the experiments are run on one NVIDIA A100 80GB GPU.
@@ Appendix B Datasets
It is a widely used benchmark for extractive question answering Rajpurkar et al. (2016). Each example consists of a question and a context paragraph from Wikipedia, with the answer being a span within the paragraph (CC BY-SA 4.0 license).
@@
It is a large-scale dataset collected from real user queries issued to the Google search engine Kwiatkowski et al. (2019). Each question is paired with a Wikipedia page. Compared to SQuAD, NQ is more challenging due to its longer contexts and more diverse question types (Apache-2.0 license).
@@
It is a multi-hop question answering dataset where answering each question requires reasoning over multiple paragraphs including both supporting facts and distractor documents (CC BY-SA 4.0 license) Yang et al. (2018).
@@
It is a variant of the NQ Longpre et al. (2021), where part of the relevant document context is swapped with similar-looking but semantically misleading content, forcing the model to rely more heavily on context utilization (MIT License).
@@
It is designed to evaluate the context-faithfulness when knowledge conflicts Bi et al. (2024). The dataset is constructed by sampling real-world facts, generating multi-hop paths, and creating counterfactual contexts by replacing entities. We use its MC (Multi-Conflicts) subset that includes multiple counterfactuals (MIT License).
@@ Appendix C Models
OLMoE-1B-7B is a fully open-source MoE developed by researchers from the Allen Institute for AI Muennighoff et al. (2024). This model consists of 16 layers where 8 out of 64 experts are activated in each, and has 7 billion parameters but activates only 1 billion parameters per input token. We use its Instruct version in this work: OLMoE-1B-7B-0924-Instruct.
@@
This is a small version of the DeepSeek-V2 model DeepSeek-AI et al. (2024). It has 27 layers where each MoE layer consists of 2 shared experts, 64 routed experts, and 6 activated experts. This configuration results in a total of 15.7 billion parameters, with 2.4 billion activated for each token. We use the parameters of its chat version: DeepSeek-V2-Lite-Chat.
@@
MiniCPM-MoE-8x2B is a MoE variant of the MiniCPM model Hu et al. (2024). It initializes using sparse upcycling, replacing MLP layers with MoE layers. With two 2 of 8 experts activated, it results in approximately 4B activated parameters. This approach significantly boosts performance across various benchmarks, while maintaining computational efficiency.
@@
Mixtral-8x7B is built upon the Mistral 7B model but incorporates 8 experts in each layer Jiang et al. (2024). This design allows each token to access a vast pool of 47B parameters, yet only 13B parameters (2 experts in each layer) are actively used during inference. We use its Instruct version, Mixtral-8x7B-Instruct-v0.1.
@@ Appendix D Evaluation by LLM-as-a-judge
EM and F1 are widely used metrics for evaluating question answering tasks. Both are standard in benchmarks such as SQuAD and NQ, ensuring comparability and providing a comprehensive assessment of model performance. Additionally, we supplement our results with LLM-as-a-judge for OLMoE-1B-7B on the NQ-Swap and ConfiQA datasets. Specifically, we employ Qwen3-32B Yang et al. (2025) to assess the correctness of the MoE models’ predictions.
@@ Appendix F Performance on Non-QA Task
We also examine RT and CEFT on other context-dependent task, e.g., summarization. We evaluate OLMoE-1B-7B and MiniCPM-MoE-8x2B on the widely used Gigaword benchmark Rush et al. (2015), where we sample 2000/500/500 examples as the train/validation/test sets, respectively.
@@ Appendix G Performance on Task Independent on the Context
We further evaluate CEFT on the MemoTrap dataset Shi et al. (2024), a benchmark specifically designed to detect whether language models memorize and regurgitate training data. Notably, MemoTrap is independent of any additional context, and therefore does not require context faithfulness.
@@ Appendix H Comparison with Other Methods for Context Utilization
We further compare our CEFT approach with Context-Faithful Prompting (CFP) Zhou et al. (2023), Context-Aware Decoding (CAD) Shi et al. (2024), ContextCite Cohen-Wang et al. (2024), and Context DPO Bi et al. (2024) on the NQ-Swap and ConfiQA datasets. Among these, CFP, CAD, and ContextCite are unsupervised methods optimizing for prompt engineering, decoding, and context pruning, respectively. In contrast, Context DPO adopts a direct preference optimization to guide LLMs to prefer context-faithful outputs.

### v3

@@ 6 Conclusion
Looking forward, we see several exciting directions: combining mechanistic interpretability techniques such as SAE DBLP:conf/acl/KangWLWTWLWFZZ25 to further unravel MoE experts; extending our methods to other forms of specialization such as reflection DBLP:conf/acl/LiDLYWWZJZ25 and reasoning DBLP:journals/corr/abs-2506-08672. We hope this work inspires further research into aspect-specific expert discovery and optimization.
@@ Acknowledgements
Prompt template used in experiments Based on the context below, output the correct answer for the following question. Context: ...Mission commander Paul and pilot Buzz Aldrin, landed the lunar module Eagle on July 20, 1969, at 20:18 UTC. Paul became the first human to step onto the lunar surface six hours after... Question: Who took the first steps on the moon in 1969?
Figure 10: Prompt template used during training and evaluation for context-dependent tasks.
@@ Appendix A Implementation Details
We fine-tune the model for 1 epoch on SQuAD, NQ, and HotpotQA datasets, and 3 epochs on NQ-Swap and ConfiQA datasets. All experiments use the AdamW DBLP:conf/iclr/LoshchilovH19/adamw optimizer with a learning rate of 5e-4, a warmup ratio of 0.1, and cosine learning rate decay. The batch size is set to 8, and the maximum sequence length is 300. All the experiments are run on one NVIDIA A100 80GB GPU.
@@ Appendix B Datasets
It is a widely used benchmark for extractive question answering DBLP:conf/emnlp/RajpurkarZLL16/squad. Each example consists of a question and a context paragraph from Wikipedia, with the answer being a span within the paragraph (CC BY-SA 4.0 license).
@@
It is a large-scale dataset collected from real user queries issued to the Google search engine DBLP:journals/tacl/KwiatkowskiPRCP19/nq. Each question is paired with a Wikipedia page. Compared to SQuAD, NQ is more challenging due to its longer contexts and more diverse question types (Apache-2.0 license).
@@
It is a multi-hop question answering dataset where answering each question requires reasoning over multiple paragraphs including both supporting facts and distractor documents (CC BY-SA 4.0 license) DBLP:conf/emnlp/Yang0ZBCSM18/hotpotqa.
@@
It is a variant of the NQ DBLP:conf/emnlp/LongprePCRD021/nq-swap, where part of the relevant document context is swapped with similar-looking but semantically misleading content, forcing the model to rely more heavily on context utilization (MIT License).
@@
It is designed to evaluate the context-faithfulness when knowledge conflicts DBLP:journals/corr/abs-2412-15280/confliqa. The dataset is constructed by sampling real-world facts, generating multi-hop paths, and creating counterfactual contexts by replacing entities. We use its MC (Multi-Conflicts) subset that includes multiple counterfactuals (MIT License).
@@ Appendix C Models
OLMoE-1B-7B is a fully open-source MoE developed by researchers from the Allen Institute for AI DBLP:journals/corr/abs-2409-02060/olmoe. This model consists of 16 layers where 8 out of 64 experts are activated in each, and has 7 billion parameters but activates only 1 billion parameters per input token. We use its Instruct version in this work: OLMoE-1B-7B-0924-Instruct.
@@
This is a small version of the DeepSeek-V2 model DBLP:journals/corr/abs-2405-04434/deepseek-v2. It has 27 layers where each MoE layer consists of 2 shared experts, 64 routed experts, and 6 activated experts. This configuration results in a total of 15.7 billion parameters, with 2.4 billion activated for each token. We use the parameters of its chat version: DeepSeek-V2-Lite-Chat.
@@
MiniCPM-MoE-8x2B is a MoE variant of the MiniCPM model DBLP:journals/corr/abs-2404-06395/minicpm. It initializes using sparse upcycling, replacing MLP layers with MoE layers. With two 2 of 8 experts activated, it results in approximately 4B activated parameters. This approach significantly boosts performance across various benchmarks, while maintaining computational efficiency.
@@
Mixtral-8x7B is built upon the Mistral 7B model but incorporates 8 experts in each layer DBLP:journals/corr/abs-2401-04088/mixtral. This design allows each token to access a vast pool of 47B parameters, yet only 13B parameters (2 experts in each layer) are actively used during inference. We use its Instruct version, Mixtral-8x7B-Instruct-v0.1.
@@ Appendix D Evaluation by LLM-as-a-judge
EM and F1 are widely used metrics for evaluating question answering tasks. Both are standard in benchmarks such as SQuAD and NQ, ensuring comparability and providing a comprehensive assessment of model performance. Additionally, we supplement our results with LLM-as-a-judge for OLMoE-1B-7B on the NQ-Swap and ConfiQA datasets. Specifically, we employ Qwen3-32B DBLP:journals/corr/abs-2505-09388 to assess the correctness of the MoE models’ predictions.
@@ Appendix F Performance on Non-QA Task
We also examine RT and CEFT on other context-dependent task, e.g., summarization. We evaluate OLMoE-1B-7B and MiniCPM-MoE-8x2B on the widely used Gigaword benchmark DBLP:conf/emnlp/RushCW15, where we sample 2000/500/500 examples as the train/validation/test sets, respectively.
@@ Appendix G Performance on Task Independent on the Context
We further evaluate CEFT on the MemoTrap dataset DBLP:conf/naacl/ShiHLTZY24, a benchmark specifically designed to detect whether language models memorize and regurgitate training data. Notably, MemoTrap is independent of any additional context, and therefore does not require context faithfulness.
@@ Appendix H Comparison with Other Methods for Context Utilization
We further compare our CEFT approach with Context-Faithful Prompting (CFP) DBLP:conf/emnlp/ZhouZPC23, Context-Aware Decoding (CAD) DBLP:conf/naacl/ShiHLTZY24, ContextCite DBLP:conf/nips/Cohen-WangSGM24, and Context DPO DBLP:journals/corr/abs-2412-15280/confliqa on the NQ-Swap and ConfiQA datasets. Among these, CFP, CAD, and ContextCite are unsupervised methods optimizing for prompt engineering, decoding, and context pruning, respectively. In contrast, Context DPO adopts a direct preference optimization to guide LLMs to prefer context-faithful outputs.